PRINTER VERSION BACK NEX Chapter 22, Problem 052 An electron enters a region of uniform electric field with an initial velocity of 57 km/s in the same direction as the electric fleld, which has magnitude E- 52 N/C. (a) What is the speed of the electron 1.9 ns after entering this region? (b) How far does the electron travel during the 1.9 ns interval? (a) Number (b) Number Units Units SHOW HINT
Answer #1

E =52 N/C

VO =57Km/sec = 57000m/sec

q= 1.6*10-19 C

F= q*E =83.2*10-19 N

m =9.1*10-31Kg

a =F/m = 9.14*1012 m/sec2

t= 1.9 ns =1.9*10-9

a) V = V0 +at =57000 + 9.14*1012 *1.9*10-9 =74366 m/sec =74.3Km/sec

b) S =V0*t +0.5*a*t2 =124.79*10-6 m =124.7 micro m